OpenStudy (anonymous):

Which of the following is a solution of x2 + 5x = -2

OpenStudy (anonymous):

is is x^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well you have to equate to zero

OpenStudy (anonymous):

then solve

OpenStudy (anonymous):

x^2+5x+2=0

OpenStudy (anonymous):

\[\frac{ 5\pm \sqrt{33} }{ 2 }\] \[\frac{ 5\pm \sqrt{17} }{ 2 }\] \[\frac{ -5\pm \sqrt{33} }{ 2 }\] \[\frac{ -5\pm \sqrt{17} }{ 2 }\]

OpenStudy (anonymous):

Is it the first one?

OpenStudy (anonymous):

yeah -bplus or minus square root b^2-4ac divide by 2(a)

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

how did the 5 become negative?

OpenStudy (anonymous):

in the formula its -b so its -5

OpenStudy (anonymous):

oh, alright. sorry I missed that. thanks you!

OpenStudy (anonymous):

no prob :)